OpenStudy (anonymous):

Please help! Will give medal. A rectangular corral is to be built with a required area, k. If an existing fence is to be used as one of the sides, determine the dimension of the corral which would cost the least.

OpenStudy (anonymous):

l = length w = width k = area

OpenStudy (igreen):

Any more information? @danielle02

OpenStudy (anonymous):

@iGreen That's all I got. I have to solve this using calculus but I don't know how to start.

OpenStudy (perl):



OpenStudy (perl):

we are given that x*y = k and lets say it costs c dollars per foot of fence total cost = c*2x + c*y

OpenStudy (anonymous):

@perl So how do I find the dimensions?

OpenStudy (perl):

you can factor out c : Cost = c*(2x + y)

OpenStudy (perl):

cost (x) = c * ( 2x + y) we are given that x*y = k, so y = k / x substitute cost(x) = c * ( 2x + k/x )

OpenStudy (perl):

we can minimize that function , since it is in one unknown

OpenStudy (perl):

find the derivative of that

OpenStudy (anonymous):

how did it become f'(x)=c[2-(k/x^2)] ?

OpenStudy (perl):

$$ \Large f(x) = c * ( 2x + \frac{k}{x} ) \\ \Large f(x) = c * ( 2x + k \cdot x^{-1} ) \\ \Large\color{red}{\Large f~'(x) = c * ( 2 \cdot 1 + k \cdot (-1) x^{-2} ) }\\ \Large f ~' (x) = c * ( 2 - \frac{k}{x^2})\\ \Large f ' (x) = 0 \iff c * ( 2 - \frac{k}{x^2}) = 0 \\ \text {divide both sides by c}\\ \Large 2 - \frac{k}{x^2} =0\\ \Large 2 = \frac{k}{x^2}\\ \Large 2x^2 = k \\ \Large x^2 = \frac{k}{2} \\ \Large x = \pm \sqrt{\frac{k}{2}} \\ \text {since length is positive}\\ \Large x= \sqrt{\frac{k}{2}} $$

OpenStudy (anonymous):

So that would be the dimension? Or do I still need to work on another equation?

OpenStudy (perl):

in the drawing how would you define length?

OpenStudy (perl):



OpenStudy (perl):

we found the dimension that minimizes cost is x = sqrt(k/2) is this width?

OpenStudy (perl):

to me , length and width are ambiguous

OpenStudy (anonymous):

@perl Thank you for helping me understand the problem. I appreciate all your effort. Thanks again!

OpenStudy (anonymous):

There was a hint that the width= twice the length

OpenStudy (perl):

yes , do you see why?

OpenStudy (perl):

$$ \text {we found that}\\ \Large x= \sqrt{\frac{k}{2}} = \frac{\sqrt{k}}{\sqrt{2}} = \frac{\sqrt{k}} {\sqrt{2}}\cdot\color{red}{\frac{\sqrt{2}}{\sqrt{2}}}= \frac{\sqrt{2}\cdot \sqrt{k}}{2} \\ \Large \text {but} \\ \Large x \cdot y = k \\ \Large (\frac{\sqrt{2}\cdot \sqrt{k}}{2}) \cdot y =k \\ \Large \text{ so } \\ \Large y = k \cdot (\frac{2}{\sqrt{2}\cdot \sqrt{k}}) = \frac{2k}{\sqrt{2k}}\cdot \color{red}{\frac{\sqrt{2k}}{\sqrt{2k}}}= \frac{2k\cdot \sqrt{2k}}{2k}\\ \\ \Large = \sqrt{ 2k} = \sqrt{2}\cdot \sqrt{k} $$

OpenStudy (perl):

now you see x is half of y (or if we switched the roles of x and y, then y is half of x)